# Cutting a Pancake with an Exotic Knife

### v1

@@ Abstract
In the first chapter of their classic book Concrete Mathematics, Graham, Knuth, and Patashnik consider the maximum number of pieces that can be obtained from a pancake by making $n$ cuts with a knife blade that is straight, or bent into a V, or bent twice into a $\sf Z$}. We extend their work by considering knives, or “cookie-cutters”, of even more exotic shapes, including a $k$-armed V, a chain of $k$ connected line segments, a long-legged version of one of the letters $\sf A$}, $\sf E$}, $\sf L$}, $\sf M$}, $\sf T$}, $\sf W$}, or $\sf X$}, a convex polygon, a circle, a figure 8, a pentagram, a hexagram, or a lollipop. In many cases a counting argument combined with Euler’s formula produces an explicit expression for the maximum number of pieces. “Strict” versions of the letters $\sf A$} and $\sf T$} are also considered, for which we have only conjectural formulas.
@@ 1 Introduction
If $S$ is an infinite straight line (or knife) $\sf K$}, we get the classical pancake-cutting problem [2, pp. 72–73], [3, Ch. 20], [4, §1.2], [12, p. 13, Problem 44], for which the answer is that the maximum number of pieces of an infinite pancake that can be obtained with $n$ cuts is
@@
The sequences, associated shapes $S$, and section numbers if they are mentioned in this article are: A000124 ($S$ is a line, §2), A000125 (plane in three dimensions), A046127 (sphere in three dimensions), A051890 (ellipse), A058331 (hyperbola), A069894 (square or convex quadrilateral, §8.1), A077588 (triangle, §1), A077591 (concave quadrilateral), A080856 (4-chain, §5), A117625 ($\sf Z$} or zig-zag, §7), A125201 ($\sf W$} or $\sf M$}, §7), A130883 (V, §4), A140064 ($\sf A$}, $\sf E$}, 3-armed V, 3-chain, Wu ( ), or Shah ( ), §§4, 6), A143689 (constrained $\overline{{\sf L}}$, §8.1), A330451 (constrained $\overline{\sf A}$, §9.2), A383464 (4-armed V, §4), A383465 (5-chain, §5), A383466 (pentagram, §10.5), A383477 (hexagram, §10.5), A386480 (circle, §10.3), A386485 (pentagon, §10.1), A386486 (8, §10.4), A389608 (lollipop ( ), §11), and A389614 (constrained $\overline{\sf T}$, §9.1), In a few cases, where we do not know the exact number of regions, the sequence mentioned is the best upper bound presently known. The second and third entries in the above list, A000125 and A046127, are the only three-dimensional problems.
@@
are discussed in the first chapter of Graham, Knuth, and Patashnik’s Concrete Mathematics [4]. In order to generalize their examples, we observe that they may be regarded as “long-legged” versions of the upper-case letters $\sf I$}, V, and $\sf Z$}. (The term was suggested by W. B. Yeats’s poem in praise of silence, Long-Legged Fly [13, pp. 327–328].)
This led us to consider long-legged versions of other letters. Informally, a long-legged version of a letter is obtained by replacing any protruding arms or legs by (semi-infinite) rays, indicated by arrow-heads, and allowing arbitrary angles between these rays. For example there is no restriction on the angle between the two arms of a long-legged V. We will, of course, be careful to give a precise definition for each shape. There are interesting long-legged versions of ${\sf A},{\sf E},{\sf L},{\sf M},{\sf T},{\sf W}$, and $\sf X$}, sometimes more than one. Curved letters (${\sf B},{\sf C},{\sf D},{\sf G},{\sf P},{\sf Q},{\sf R},{\sf S}$, $\ldots$) do not seem appropriate for this treatment, since as we shall show below, once curved lines are permitted (without any bound on their curvature), it is easy to create arbitrarily many regions. The curved shapes $\sf O$}10.3), 810.4) and the lollipop (§11) are exceptions. A cookie-cutter in the shape of a long-legged $\sf A$} initially looked especially challenging, although as we shall see in §6, it has a surprisingly simple solution. The constrained long-legged $\bar{\sf A}$, however (§9.2), appears to be genuinely hard.
@@
These constrained letters appear to be more difficult to analyze, and we will therefore divide the shapes we consider into two families. Members of what we will call the affine family of shapes have the property that if $S$ is a shape of that type, so is any image of $S$ under the full two-dimensional affine group ${\mathcal{G}}_{\mathnormal{a}ff}$ generated by linear transformations (which need not be isometries), followed by a translation. On the other hand, the similarity family of shapes have the property that if $S$ is a shape of that type, so is any image of $S$ under the group ${\mathcal{G}}_{\mathnormal{s}im}$ of two-dimensional similarities, generated by isometries (rotations, reflections, translations), followed by a dilation. The affine shapes we consider are the line, the hatpin, the $k$-armed V, the $k$-chain, the long-legged letters $\sf A$}, $\sf M$}, $\sf W$}, $\sf Z$} (§§2-7), and the convex polygon (§10.1). All the other shapes belong to the similarity family.
@@
Subsequent sections are devoted to a $k$-chain (§5), and long-legged versions of $\sf A$}6), $\sf Z$}, $\sf W$}, and $\sf M$}7), and the constrained long-legged letters $\bar{{\sf L}}$ and $\bar{{\sf X}}$8), and $\bar{{\sf T}}$ and $\bar{{\sf A}}$9). We study several finite shapes in §10, namely a convex $k$-sided polygon (§10.1, §10.2), a circle (§10.3), a figure 810.4), and a regular pentagram and hexagram (§10.5). Our final shape is a lollipop ( ) (§11).
@@
Notation. In order to distinguish between the letters themselves and the symbols denoting their properties, we will use a sans serif font (A, L, V, etc.) for the letters and other shapes. We use K for the infinite knife-cuts in the basic pancake-cutting problem, H for the hatpin, $k{\sf V}$ for the $k$-armed V, $k{\sf C}$ for the $k$-chain, and $k{\sf P}$ for the $k$-sided polygon. An overbar ($\bar{{\sf L}}$, $\bar{{\sf X}}$, etc.) indicates a constrained shape.
@@
The parameters of $G_{S}(n)$ will be denoted by $V_{{\mathcal{B}}}$, $_{{\mathcal{C}}}, V=V$$V_{{\mathcal{B}}}+V_{{\mathcal{C}}}$ (numbers of base nodes, crossing nodes, total vertices), $E_{\infty}$}, $_{f}, E=E$${E_{\infty}}+E_{f}$ (infinite edges, finite edges, total edges), and $R_{\infty}$}, $_{f}, R=R$${R_{\infty}}+R_{f}$ (infinite regions, finite regions, total regions). These definitions will be repeated when they appear below.
@@ 2 The planar graph and the maximum number of regions
We were surprised not to find (3) in Lakatos [8], a book devoted to proofs of different versions of Euler’s formula, nor in the standard books on finite graph theory. But (3) is simple to prove. Draw a rectangle around $G_{S}(n)$, large enough so that only the $E_{\infty}$} edges cross it. The rectangle therefore has $4+{E_{\infty}}$ edges and $4+{E_{\infty}}$ vertices. The graph formed by the union of $G_{S}(n)$ and the rectangle has $E+4+{E_{\infty}}$ edges, $V+4+{E_{\infty}}$ vertices, and by (2) has $(E+4+{E_{\infty}})-(V+4+{E_{\infty}})+1=V-E+1$ regions inside the rectangle.
@@
In most of the examples studied in this paper, we are able to achieve equality in (11). The chief difficulty, once we have determined $\kappa(S)$, lies in placing the $n$ copies of $S$ so that every pair intersect in $\kappa(S)$ points. We run into difficulties if, once one copy of $S$ has been positioned, there are only a small range of possibilities for the second (and therefore all) subsequent copies. This is the case for the constrained letters $\overline{\sf T}$9.1) and $\overline{\sf A}$9.2), and for the figure $\sf 8$}10.4) and the lollipop (§11).
This difficulty seems never to arise with any of our affine family of shapes, where there always seems to be a plentiful supply of candidates for the next copy of the shape. This leads us to make the following bold
@@ 5 The $k$-chain
Figure 8: Transforming a pancake graph $G_{\sf K}(3)$ to an optimal $G_{3{\sf C}}(3)$ graph with $a_{3{\sf C}}(3)=34$ regions, by replacing hatpins by long narrow 3-chains.
@@ 6 The long-legged A
Our drawings suggested that the initial values of $a_{\sf A}(n)$, the maximum numbers of regions achievable with $n$ $\sf A$}’s, for $n=1,2$, and 3, were 3, 14 (Fig. 11), and 34, matching the counts for 3-armed Vs and later, the counts for 3-chains. This was not a coincidence!
@@
Sketch of proof. We have already established the theorem for 3-armed V’s (in §4), and for 3-chains (in §5). It remains to consider long-legged $\sf A$}’s.
@@
On the other hand, to convert the long-legged $\sf A$} graph $G$ into an arrangement of 33-chains, we simply apply the transformation shown in Fig. 13. This preserves the number of regions (we invite the reader to apply it to Fig. 11 and check that the result still has 14 regions).
@@ 7 The long-legged Z, W, and M
8 The constrained long-legged $\overline{{\sf L}}$ and $\overline{{\sf X}}$
In this section we discuss our first two shapes from the similarity family, $\overline{\sf L}$ and $\overline{{\sf X}}$, both of which we are able to solve.
@@ 8.1 The constrained long-legged $\overline{\sf L}$
With no restriction on the angle, a long-legged $\sf L$} would be indistinguishable from a long-legged V. We define a constrained long-legged $\bar{{\sf L}}$ to consist of a distinguished point, its center (e.g., $(0,0)$), from which two perpendicular rays emanate (e.g., $\{y=0,x\geq 0\}$ and $\{x=0,y\geq 0\}$), or any of the images of such a figure under the similarity group ${\mathcal{G}}_{\mathnormal{s}im}$.
Two $\bar{{\sf L}}$’s can intersect in at most $\kappa({\bar{\sf L}})=3$ points, so $V_{{\mathcal{C}}}\leq 3\binom{n}{2}$ and then (6) gives
@@
Figure 15: The construction of $n$ constrained long-legged $\bar{{\sf L}}$’s, showing the case $n=5$. There are $36$ regions. (The arrowheads have been omitted.)
Equality in (17) can be achieved by modifying the standard construction for drawing $n$ squares with the maximum number of regions (see A069894). This construction starts by marking $4n$ equally-spaced points $P_{0},\ldots,P_{4n-1}$ around the circumference of a circle. To get an optimal arrangement of $n$ squares, we would take the squares with vertices $\{P_{i},P_{i+n},P_{i+2n},P_{i+3n}\}$ for $i=0,1,\ldots,n-1$. To get an optimal arrangement of $n$ $\bar{{\sf L}}$’s, we take the $\bar{{\sf L}}$’s formed by the top left corners of these squares. That is, we draw an $\bar{{\sf L}}$ centered at $P_{i}$ with arms $P_{i}-P_{i+n}$ and $P_{i}-P_{i+3n}$, for $i=0,\ldots,n-1$ . Figure 15 shows the construction when $n=5$. In general the graph contains $2n$ infinite regions, there are $n$ rows of n1n-1 cells each, and a further triangle of (n1)(n2)/2(n-1)(n-2)/2 cells, for a total of $R=a_{\bar{{\sf L}}}(n)=(3n^{2}-n+2)/2$ regions. For $n\geq 0$ the values are $1,2,6,13,36,52,71,\ldots$ (A143689). The OEIS entry makes no mention of this geometric application, so this result may be new.
@@ 8.2 The constrained long-legged $\overline{\sf X}$
We can indeed achieve equality in (18) by modifying the construction of pancake graphs given in §2. To construct a $G_{\overline{\sf X}}(n)$ containing $n$ copies of $\overline{\sf X}$, set $\theta=\pi/(4n)$, and let $\lambda_{t}$ denote the line (9). Draw the lines $\lambda_{t}$ for $t=0,1,\ldots,3n-1$ (now we need more of these lines than when we constructed the pancake graph). The $t$th copy of $\sf X$} is formed from the lines $\lambda_{t}$ and $\lambda_{t+2n}$ for $t=0,\ldots,n-1$. Figure 16 illustrates the construction for $n=3$. Of all our constructions, this one seems to be the most likely to have applications, possibly to some abstract version of the Gerrymandering problem.

### v2

@@ Abstract
In the first chapter of their classic book Concrete Mathematics, Graham, Knuth, and Patashnik consider the maximum number of pieces that can be obtained from a pancake by making $n$ cuts with a knife blade that is straight, or bent into a V, or bent twice into a $\sf Z$}. We extend their work by considering knives, or “cookie-cutters”, of even more exotic shapes, including a $k$-armed V, a chain of $k$ connected line segments, a long-legged version of one of the letters $\sf A$}, $\sf E$}, $\sf H$}, $\sf L$}, $\sf M$}, $\sf T$}, $\sf W$}, or $\sf X$}, a convex polygon, a circle, a $\phi$}, a figure 8, a pentagram, a hexagram, or a lollipop. In many cases a counting argument combined with Euler’s formula produces an explicit expression for the maximum number of pieces. “Constrained” versions of the long-legged letters $\sf A$} and $\sf T$} are also considered, for which we have only conjectural formulas.
@@ 1 Introduction
If $S$ is an infinite straight line (or knife) $\sf K$}, we get the classical pancake-cutting problem [2, pp. 72–73], [3, Ch. 20], [4, §1.2], [14, p. 13, Problem 44], for which the answer is that the maximum number of pieces of an infinite pancake that can be obtained with $n$ cuts is
@@
The sequences, associated shapes $S$, and section numbers if they are mentioned in this article are: A000124 ($S$ is a line, §2), A000125 (plane in three dimensions), A046127 (sphere in three dimensions), A051890 (ellipse), A058331 (hyperbola), A069894 (square or convex quadrilateral, §8.1), A077588 (triangle, §1), A077591 (concave quadrilateral), A080856 (4-chain, §5), A117625 ($\sf Z$} or zig-zag, §7), A125201 ($\sf W$} or $\sf M$}, §7), A130883 (V, §4), A140063 (constrained $\overline{\sf H}$ or $\overline{\phi}$, §8.3), A140064 ($\sf A$}, $\sf E$}, 3-armed V, 3-chain, Wu ( ), or Shah ( ), §§4, 6), A143689 (constrained $\overline{{\sf L}}$, §8.1), A330451 (constrained $\overline{\sf A}$, §9.2), A383464 (4-armed V, §4), A383465 (5-chain or H, §§5, 8.3), A383466 (pentagram, §10.5), A386477 (hexagram, §10.5), A386480 (circle, §10.3), A386485 (pentagon, §10.1), A386486 (8, §10.4), A389608 (lollipop ( ), §11), and A389614 (constrained $\overline{\sf T}$, §9.1), In a few cases, where we do not know the exact number of regions, the sequence mentioned is the best upper bound presently known. The second and third entries in the above list, A000125 and A046127, are the only three-dimensional problems.
@@
are discussed in the first chapter of Graham, Knuth, and Patashnik’s Concrete Mathematics [4]. In order to generalize their examples, we observe that they may be regarded as “long-legged” versions of the upper-case letters $\sf I$}, V, and $\sf Z$}. (The term was suggested by W. B. Yeats’s poem in praise of silence, Long-Legged Fly [15, pp. 327–328].)
This led us to consider long-legged versions of other letters. Informally, a long-legged version of a letter is obtained by replacing any protruding arms or legs by (semi-infinite) rays, indicated by arrow-heads, and allowing arbitrary angles between these rays. For example there is no restriction on the angle between the two arms of a long-legged V. We will, of course, be careful to give a precise definition for each shape. There are interesting long-legged versions of ${\sf A},{\sf E},{\sf H},{\sf L},{\sf M},{\sf T},{\sf W}$, and $\sf X$}, sometimes more than one. Curved letters (${\sf B},{\sf C},{\sf D},{\sf G},{\sf P},{\sf Q},{\sf R},{\sf S}$, $\ldots$) do not seem appropriate for this treatment, since as we shall show below, once curved lines are permitted (without any bound on their curvature), it is easy to create arbitrarily many regions. The curved shapes $\overline{\phi}$8.3), $\sf O$}10.3), 810.4) and the lollipop (§11) are exceptions. A cookie-cutter in the shape of a long-legged $\sf A$} initially looked especially challenging, although as we shall see in §6, it has a surprisingly simple solution. The constrained long-legged $\overline{\sf A}$, however (§9.2), appears to be genuinely hard.
@@
These constrained letters appear to be more difficult to analyze, and we will therefore divide the shapes we consider into two families. Members of what we will call the affine family of shapes have the property that if $S$ is a shape of that type, so is any image of $S$ under the full two-dimensional affine group ${\mathcal{G}}_{\mathnormal{a}ff}$ generated by linear transformations (which need not be isometries), followed by a translation. On the other hand, the similarity family of shapes have the property that if $S$ is a shape of that type, so is any image of $S$ under the group ${\mathcal{G}}_{\mathnormal{s}im}$ of two-dimensional similarities, generated by isometries (rotations, reflections, translations), followed by a dilation. The affine shapes we consider are the line, the hatpin, the $k$-armed V, the $k$-chain, the long-legged letters $\sf A$}, $\sf H$}, $\sf M$}, $\sf W$}, $\sf Z$} (§§2-7, 8.3), and the convex polygon (§10.1). All the other shapes belong to the similarity family.
@@
Subsequent sections are devoted to a $k$-chain (§5), and long-legged versions of $\sf A$}6), $\sf Z$}, $\sf W$}, and $\sf M$}7), and the constrained long-legged letters $\overline{{\sf L}}$, $\overline{{\sf X}}$, $\overline{\sf H}$8), $\overline{{\sf T}}$, and $\overline{{\sf A}}$9). We study several finite shapes in §10, namely a convex $k$-sided polygon (§10.1, §10.2), a circle (§10.3), a figure 810.4), and a regular pentagram and hexagram (§10.5). Our final shape is a lollipop ( ) (§11).
@@
Notation. In order to distinguish between the letters themselves and the symbols denoting their properties, we will use a sans serif font (A, L, V, etc.) for the letters and other shapes. We use K for the infinite knife-cuts in the basic pancake-cutting problem, H for the hatpin, $k{\sf V}$ for the $k$-armed V, $k{\sf C}$ for the $k$-chain, and $k{\sf P}$ for the $k$-sided polygon. An overbar ($\overline{{\sf L}}$, $\overline{{\sf X}}$, etc.) indicates a constrained shape.
@@
The parameters of $G_{S}(n)$ will be denoted by $V_{{\mathcal{B}}}$, $_{{\mathcal{C}}}, V=V$$V_{{\mathcal{B}}}+V_{{\mathcal{C}}}$ (numbers of base nodes, crossing nodes, total vertices), $E_{\infty}$}, $_{f}, E=E$${E_{\infty}}+E_{f}$ (infinite edges, finite edges, total edges), and $R_{\infty}$}, $_{f}, R=R$${R_{\infty}}+R_{f}$ (infinite regions, finite regions, total regions). These definitions will be repeated when the symbols first appear below.
@@ 2 The planar graph and the maximum number of regions
We were surprised not to find (3) in Lakatos [9], a book devoted to proofs of different versions of Euler’s formula, nor in the standard books on finite graph theory. But (3) is simple to prove. Draw a rectangle around $G_{S}(n)$, large enough so that only the $E_{\infty}$} edges cross it. The rectangle therefore has $4+{E_{\infty}}$ edges and $4+{E_{\infty}}$ vertices. The graph formed by the intersection of $G_{S}(n)$ and the rectangle has $E+4+{E_{\infty}}$ edges, $V+4+{E_{\infty}}$ vertices, and by (2) has $(E+4+{E_{\infty}})-(V+4+{E_{\infty}})+1=V-E+1$ regions inside the rectangle.
@@
In most of the examples studied in this paper, we are able to achieve equality in (11). The chief difficulty, once we have determined $\kappa(S)$, lies in placing the $n$ copies of $S$ so that every pair intersect in $\kappa(S)$ points. We run into difficulties if, once one copy of $S$ has been positioned, there is only a small range of possibilities for the second (and therefore all) subsequent copies. This is the case for the constrained letters $\overline{\sf T}$9.1) and $\overline{\sf A}$9.2), the figure $\sf 8$}10.4), and the lollipop (§11).
This difficulty seems not to arise with any of our affine family of shapes, where there always seems to be a plentiful supply of candidates for the next copy of the shape. This leads us to make the following bold
@@ 5 The $k$-chain
Figure 8: Transforming a pancake graph $G_{\sf K}(3)$ to an optimal $G_{3{\sf C}}(3)$ graph with $a_{3{\sf C}}(3)=34$ regions, by replacing the knife cuts by long narrow 3-chains.
@@ 6 The long-legged A
Our drawings suggested that the initial values of $a_{\sf A}(n)$, the maximum numbers of regions achievable with $n$ $\sf A$}’s, for $n=1,2$, and 3, were 3, 14 (Fig. 11), and 34, matching the counts for both 3-armed Vs and 3-chains. This was not a coincidence!
@@
Sketch of proof. We have already established the theorem for 3-armed V’s (in §4) and 3-chains (in §5). It remains to consider long-legged $\sf A$}’s.
@@
On the other hand, to convert the long-legged $\sf A$} graph $G$ into an arrangement of 33-chains, we simply apply the transformation shown in Fig. 13 to each $\sf A$}. This preserves the number of regions (we invite the reader to apply it to Fig. 11 and check that the result still has 14 regions).
@@ 7 The long-legged Z, W, and M
8 The constrained long-legged $\overline{{\sf L}}$, $\overline{{\sf X}}$, $\overline{{\sf H}}$, and $\overline{{\phi}}$
In this section we discuss our first four shapes from the similarity family, $\overline{\sf L}$, $\overline{{\sf X}}$, $\overline{{\sf X}}$, and $\overline{{\phi}}$, all of which we are able to solve.
@@ 8.1 The constrained long-legged $\overline{\sf L}$
With no restriction on the angle, a long-legged $\sf L$} would be indistinguishable from a long-legged V. We define a constrained long-legged $\overline{{\sf L}}$ to consist of a distinguished point, its center (e.g., $(0,0)$), from which two perpendicular rays emanate (e.g., $\{y=0,x\geq 0\}$ and $\{x=0,y\geq 0\}$), or any of the images of such a figure under the similarity group ${\mathcal{G}}_{\mathnormal{s}im}$.
Two $\overline{{\sf L}}$’s can intersect in at most $\kappa({\overline{\sf L}})=3$ points, so $V_{{\mathcal{C}}}\leq 3\binom{n}{2}$ and then (6) gives
@@
Figure 15: The construction of $n$ constrained long-legged $\overline{{\sf L}}$’s, showing the case $n=5$. There are $36$ regions. (The arrowheads have been omitted.)
Equality in (17) can be achieved by modifying the standard construction for drawing $n$ squares with the maximum number of regions (see A069894). This construction starts by marking $4n$ equally-spaced points $P_{0},\ldots,P_{4n-1}$ around the full circumference of a circle. To get an optimal arrangement of $n$ squares, we would take the squares with vertices $\{P_{i},P_{i+n},P_{i+2n},P_{i+3n}\}$ for $i=0,1,\ldots,n-1$. To get an optimal arrangement of $n$ $\overline{{\sf L}}$’s, we take the $\overline{{\sf L}}$’s formed by the top left corners of these squares. That is, we draw an $\overline{{\sf L}}$ centered at $P_{i}$ with arms $P_{i}-P_{i+n}$ and $P_{i}-P_{i+3n}$, for $i=0,\ldots,n-1$ . Figure 15 shows the construction when $n=5$. In general the graph contains $2n$ infinite regions, there are $n$ rows of n1n-1 cells each, and a further triangle of (n1)(n2)/2(n-1)(n-2)/2 cells, for a total of $R=a_{\overline{{\sf L}}}(n)=(3n^{2}-n+2)/2$ regions. For $n\geq 0$ the values are $1,2,6,13,36,52,71,\ldots$ (A143689). The OEIS entry makes no mention of this geometric application, so this result may be new.
@@ 8.2 The constrained long-legged $\overline{\sf X}$
We can indeed achieve equality in (18) by modifying the construction of pancake graphs given in §2. To construct a $G_{\overline{\sf X}}(n)$ containing $n$ copies of $\overline{\sf X}$, set $\theta=\pi/(4n)$, and let $\lambda_{t}$ denote the line (9). Draw the lines $\lambda_{t}$ for $t=0,1,\ldots,3n-1$ (now we need more of these lines than when we constructed the pancake graph). The $t$-th copy of $\sf X$} is formed from the lines $\lambda_{t}$ and $\lambda_{t+2n}$ for $t=0,\ldots,n-1$. Figure 16 illustrates the construction for $n=3$. Of all our constructions, this one seems to be the most likely to have applications, possibly to some abstract version of the Gerrymandering problem.
Figure 17: An unconstrained long-legged H is the same as two V’s joined by a line segment, which can be rearranged to have four self-crossings, and then (invertibly) transformed into a 5-chain by “raising the bar” (the heavy line).
8.3 The constrained long-legged $\overline{\sf H}$ and $\overline{\phi}$
Our rules for long-legged letters tell us that an (unconstrained) long-legged H is the same as two long legged V’s with their tips joined by a line segment (see Fig. 17). Then $\sigma({\sf H})=4$, $\kappa({\sf H})=25$, and with $n$ H’s, (5) tells us that $a_{\sf H}(n)\leq(25n^{2}-11n+2)/2$. This matches the counts for a $5$-chain, (16) (A383465), and equality is easily achieved since the invertible transformation shown in Fig. 17 maps an optimal graph with $n$ H’s into one with $n$ 55-chains.
Figure 18: One of the three inequivalent ways in which two constrained $\overline{\sf H}$’s can interesct in seven points.
Figure 19: Construction of $n$ $\overline{\sf H}$’s with each pair intersecting in 7 points (here $n=3$).
We can define a constrained long-legged $\overline{\sf H}$ to consist of a pair of parallel lines joined by a perpendicular line segment (the crossbar), or any image of such a figure under ${\mathcal{G}}_{\mathnormal{s}im}$. Two $\overline{\sf H}$’s can meet in a maximum of $\kappa(\overline{\sf H})=7$ points. This can be done in three inequivalent ways, but we will only make use of one of them, the arrangement shown in Fig. 18.
It is straightforward to arrange $n$ $\overline{\sf H}$’s so that each pair intersect in seven points in this way, by first drawing the $n$ crossbars so that each one crosses the others at a small angle, similar to the arrangement of lines in Fig. 4. By making these crossbars have slowly increasing and incommensurable lengths, we ensure that when we add the parallel lines to the crossbars, there are no triple intersections in the resulting graph. Figure 19 illustrates the construction when $n=3$.
From (5), the maximum number of regions with $n$ $\overline{\sf H}$’s is then
[…]
with initial values $1,4,14,31,55,\ldots$. We were surprised to find that this sequence was already in the OEIS, in a 2008 entry. This entry formerly contained a link to a webpage [8] of David Kinsella, where it was stated that this sequence gives the maximum number of regions that the plane can be divided into by drawing $n$ circles and $n$ lines. This webpage now seems to be lost, but the Internet Archive’s Wayback Machine has preserved enough of this document (see [8]: the figures are missing) to make it clear that this was an empirical observation based on examination of the constructions for $n\leq 4$ pairs of circles and lines.
Figure 20: An invertible map from $\overline{\sf H}$’s to $\overline{\phi}$’s.
A weaker version of this claim can be obtained from our analysis of the $\overline{\sf H}$ shape. Define a constrained $\overline{\phi}$ to consist of a circle with a line through its center. Two such $\overline{\phi}$’s can intersect in a maximum of seven points. An optimal graph containing $n$ $\overline{\phi}$’s, each intersecting all of the others in seven points, can be obtained from an optimal graph with $n$ $\overline{\sf H}$’s by applying the invertible transformation shown in Fig. 20. It follows that (19) applies also to $\overline{\phi}$’s. It is plausible that to obtain the maximum number of regions from $n$ circles and $n$ lines one should first pair them up into $\overline{\phi}$’s, so this is a reasonable replacement for Kinsella’s claim.